Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day08


Crew 102 Log Book for April 2, 2011

Photos of the day





Commander's Check-In Report

Franco Carbognani Reporting

Crew Physical Status: good exept for one crewmember whose cold is slowly improving

Time Departed/Returned from EVA: 12:10 - 14:00, 17:00 - 20:00

Brief Narrative of Field Mission Results:

Performed tests on soil permeability

Discovered and recorded GPS of sites abundant in cyanobacteria samples

EVA Data/Interpretations:

Rock Samples analysis ongoing

EVA GPS tracing using crew smart-phones is working well

Engineering/Hab Maintenance:

Main generator running fine after provisional fixes

Water Heater leaking problem manageable keeping heating period to the minumum.

Plans for Tomorrow:

Long EVA with ATV for localisation of Gryphea Arcuata bioerosion samples sites

Short EVA around the Hab dedicated to corrosion inspection and sample collection for optical microscopy analysis

Proceed on analysis of Radiotelescope data after calibration. Verify the possibility to define online analysis able to detect energetic solar storms.

Possibly proceeding on bringing back gray water system in operation based on the Mission Support Chief Engineer inputs.

Continue on ongoing psychological and physiological studies.

Support Requested:

Don Lusco for final fixes on power generator.


Commander's Report

Franco Carbognani Reporting

Main generator still running fine.

Two ATV EVAs has been performed. The first in order to verify EVA crew capability to find a specific spot based on GPS data, the second for site localisation and collection of Gryphea   Arcuata bioerosion samples.

Four members of the Crew went out of sim for 1 and a half hours, going to Hanksville to pick up some scientific equipment sent over for our biologic experiments and buy some badly needed consumables.

Radiotelescope data collection for off-line analysis is proceeding. Unfortunately we are missing a radiotelescope calibrator in order to get full quality data. This is supposed to arrive on-site within 2 weeks (too late....)

Hab Monitoring and Alarming System up and running in provisional mode with corresponding main web page accessible internally. Set 2 audible alarms: one on Hab temperature going over 45 C° and one on CO2 concentration going over 600ppm

Green Hab: started pumping water from the gray water tank into the straining tank in order to fill the Trickle Reservoirs.

The Telescope mount and controller are ready in the Observatory to be handed over to Don Lusco.

Mood of the crew is good.

The crewmembers cold is getting better.

On water issue side, we had the final word of Don Lusco, confirming the hypothesis of dissolved salts. So the "mystery" is explained and matter closed.


Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No):Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Kitty provisional repair seems holding. The cooling fan is running smoothly as is the whole Kitty engine.

Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.64

Propane Tank (Full/0.75/0.5/0.25/Empty): 51%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.60

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 4/4

Honey Oil Quantity (# of Quarts in storage): 0/4, to be refilled

ATV Oil Quantity (# of Quarts in storage): all 4/4

Car Oil Quantity (# of Quarts in storage): 3/4


Water Status:

Outside Potable Water Tank Level (inches from bottom): 17.2

Trailer Potable Water Tank Level (inches from bottom): 17

Hab Potable Water Tank Level (inches from bottom): 30

Potable Water Meter Reading: 44390,3

In to GreenHab Meter Reading:

Notes/Comments:

Water heater leaking problem: by limiting the heating time we have see no recurrence of the problem on the last 3 days

The Trailer Potable Water Tank is now almost less than a half filled. May be worth planning for refilling.


GreenHab:

Condition of plants in Tank 1:

Condition of plants in Tank 2:

Condition of plants in Tank 3:

Greenhab Notes/Comments:

Gray Water System: started pumping water from the gray water tank into the straining in order to fill the Trickle Reservoirs.


Observatory:

Telescope Used? (Yes/No):No

Observatory Notes/Comments:

Boxes containing the telescope mount and controller are ready in the Observatory in order to be handed over to Don Lusco.


Transportation: Hab Car (NOT RENTAL) Used (yes/no): yes

Oil Checked (yes/no):yes

Tire Status: good

Odometer at end of day: 123775.7

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used:yes

Oil Checked: n

Fuel Consumed: 1/2 liter

Tire Status: inflated

Comments:


ATV 2 (Spirit):

Used: no

Oil Checked: n

Fuel Consumed: none

Tire Status: inflated

Comments: this ATV does not start


ATV 3 (Viking I):

Used: yes

Oil Checked: n

Fuel Consumed: 1/2 liter

Tire Status: good

Comments:


ATV 4 (Viking II):

Used: yes

Oil Checked: n> Fuel Consumed: 1/2 liter

Tire Status: good

Comments:>

Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65

Actual temperature upstairs (Farenheit): 77

Thermostat setting downstairs (Farenheit):55

Actual temperature downstairs (Farenheit


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering Notes/Comments:

none


EVA 8 Report

Kavya K.Manyapu Reporting

Time of Departure: 11:00

Time of Arrival: 12.30

Crew Members: K.Manyapu, J.Deaton, A.Kapoglou

Site Location1: Dinosaur Fossils

Site Location2: Exploration

Altitude: 4505 feet

Total Distance: 5.2 miles (complete trip), 0.8 miles to the Dinosaur fossils

Transit Mode: ETV

Objectives: The main purpose of this EVA was to check the ability of the crew to re-trace the navigation path that was recorded during a previous EVA by different crew members using a GPS. The route to the Dinosaur fossils was to be tracked by the present EVA crew using just the data that was traced by the previous crew without additional support from the cap-com or crew who have been there.

Results: The crew was successful in reaching the destination and exploring the Dinosaur fossils in the area. Furthermore the crew set on an exploration mission further north from the fossil area on the ETVs about 3 miles north of the MDRS habitat.

Lessons Learned:

The GPS application from apple (iPhone/iTouch) called Everytrail is much more reliable than the GPS at MDRS and much more user friendly as well. The data can be uploaded to the computer and used for various purposes.


EVA 9 Report

Yuval Brodsky Reporting

Time: 1655 - 1934

Crew Members: Y. Brodsky (EVA CDR), L. Vimercati, A. Kapoglou

Site Location: Radio ridge

Transit Mode: ATV

Objectives: Geo-tagging of corrosion and bio-erosion of Gryphea arcuata samples on Radio ridge

Results: Numerous samples of corrosion and bio-erosion were successfully located and geo-tagged. Subsequent EVAs to Radio ridge will be undertaken in order to catalog further samples.

Lessons Learned: N/A


Chef's Report

Dr. John Deaton Reporting

Time: 22:00

1a) Today was a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

n/a

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Forever Young pasta

Kung fu chicken

Boiled vegetables

Left-over noodle soap from yesterday

Left-over Shepard's Pie filler from yesterday

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

All food was consumed

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

n/a

5) Lessons learned

Frequent stirring is important to prevent food from sticking to pan

6) Comments/questions for Kim and Jean, the food study investigators

n/a


Journalist Report

Kavya Manyapu Reporting

And the countdown to return to Earth starts. T-7 days!

The crew is busy conducting experiments and validating some of the results. Today’s first EVA was commanded by K. Manyapu accompanied by J.Deaton and A.Kapoglou. One of the goals of the EVA was to perform a navigation re-tracking task using the GPS to locate the destination that was tracked during an earlier EVA to the Dinosaur fossil location. The crew was successful in reaching the destination by using the GPS data recorded by the previous EVA. The other part of the EVA involved more exploration of the surrounding area using ETVs. Y.Brodky served as the Cap-com for the EVA. K.Manyapu and F.Carbognani took a quick tour of the engineering aspects of the hab for maintenance this afternoon.  Crew Biologist L.Vimercati started culturing the spirulina in a medium with 1:5 ratios of urine and gray water.  The second EVA of the day was commanded by Y.Brodsky in search of bio-erosion who was accompanied by L.Vimercati and A.Kapoglou. The cap-com for this EVA was K.Manyapu. The EVA lasted 2.40 hours. Samples for gryphea arcuta were collected during the EVA. The other experiments continue to take place alongside the EVAs.

That’s all for now from Crew102.